OpenStudy (anonymous):

Solve for x and y A. 3x+8y=2 B. x=2n -3x-2y=-5 y= 5n 2x+y=54

OpenStudy (anonymous):

a add : 6y = -3 y = - 1/2 sub y = 1/2 in first equation: 3x + 8(-1/2) = 2 3x = 4 + 2 x = 2

OpenStudy (anonymous):

b. plug x and y in last equn: 2(2n) + 5n = 54 9n = 54 n = 6 so x = 2n = 12 and y = 5n = 30